wiki:NDGMonthly

Version 16 (modified by pjkersha, 11 years ago) (diff)

--

10th April 2008

Agenda

  1. Current Status/ short term goals
    1. what work and support is required to get all sites running services?
      1. Discovery record supply pipeline - OK at all sites?
      2. Browse?
      3. Data available via CSML, or other methods?
      4. WMS, WCS?
      5. Security?
        • BADC - installed and running for NDG Browse. Phil is working on NDG Security interface for BADC data browser
        • BODC
        • NOCS - awaiting re-configuration of server
        • PML
          • installed and working with cross-site access: Mike can log in as a PML user and access BADC datasets. Phil can log into BADC from PML.
          • TODO: secure some PML MOLES/CSML
          • Question: way forward for securing datasets: through MOLES schema (roles+Attribute Authority address in data) or other?
  1. Way Forward
    1. Medium term initiative - Bryan. What will this actually fund? How is it organised?
    2. Any developments need to be incorporated in a manged way with:-
      1. clear set of agreed release candidates
      2. prioritised
      3. roadmap

6th February 2008

Actions

  1. Bryan to send around updated details on encoding Browse links in Moles/DIF ("VIEW EXTENDED METADATA")
  2. Bryan send around link to NERC Portals GUI
  3. Sue send around script for doing moles2dif with ndgUtils pythonic accessing of eXist.
  4. BODC - priority is getting some data available for visualisation(i.e. some CSML). Tide gauge is favourite.
  5. Give Sue/Bryan? some examples of their moles records for trying out with unit tests/Browse.
  6. BODC looking at getting browse/stack working on their production server (but changes to this server are limited to POL support, so slow process).
  7. PML priority is to push browse/stack to live.
  8. PML may look into security next week.
  9. BADC/NEODC priority is to move browse service to a production machine.
  10. BADC/NEODC also need to move security to production machine.

Agenda

  1. NERC Data Discovery Service status
  2. NERC Portals project status
  3. What is status at DPs for Pylons Browse/Security?/WMS/WCS installs?
  4. Content
    • MOLES?
    • DIFs?
    • CSML?
  5. AOB
  6.  http://proj.badc.rl.ac.uk/ndg/roadmap

13th December 2007

Actions

  1. Be available for advising Fabio tomorrow (Friday 14th Dec)
  2. Bryan send around example of URL group in DIF for a Browse link.
  3. Phil try pylons restart on NOCS server.
  4. Roy/Ray? give Sue example of existing Moles record so she can test for availability of BODC browse.
  5. Mike see if there's a latest python2.4 egg and install (for browse improvements).
  6. Sue ticket desirability of service checking crontab on Production NERC-DDS.
  7. Sue arrange next meeting.

Agenda

  1. NERC Data Discovery Service status
  2. NERC Portals project status
  3. What is status at DPs for Pylons Browse/Security?/WMS/WCS installs?
    • BADC/NEODC - currently using glue until NERC DDS done. Then will install on CEDA production environments.
    • PML - browse and security up to date working (need to create 'real' test users), csml granule issues being worked on
    • BODC - #886?
    • NOCS - ?
  4. Security - integration of MyProxy with DPs user databases for Deployment
    • up until now login has been based on credentials held in MyProxy instead of username/password in DPs user database
    • MyProxy can be configured to make a call out to a user database and query the relevant tables for username/password. For this, each site would manage its own CA
    • this is tested at BADC
    • co-ordinate roll out at other DPs ...
  5. Content
    • MOLES?
    • DIFs?
    • CSML?
  1.  http://proj.badc.rl.ac.uk/ndg/roadmap